OpenStudy (anonymous):

WILL BECOME A FAN AND A GIVEL A MEDAL Let f(x)= ln(3x+2)^k for some positive constant k. If f'(2)=3, what is the value of k? (A) ln3/ln8 (B) ln8 (C) 4 (D) 8 (E) 16

OpenStudy (anonymous):

first you know that \[\ln(3x+2)^k=k\ln(3x+2)\]

OpenStudy (anonymous):

the derivative is therefore \[\frac{3k}{3x+2}\] replace \(x\) by \(2\), set the result equal to \(3\) and solve for \(k\)

OpenStudy (anonymous):

I did that \[3=\frac{ 3k }{ 3(2)+2 } =\frac{ 3k }{ 8 }\] then I multiplied the 8 to the other side \[18 \times 3= 3k\] then I divided the 3 \[18=k\] and that is not any of the choices. Can anyone see my mistake?